# The Past, Present, and Future of Automation in Model-Driven Engineering

### v1

@@ 2.2.1. Formal Methods
Due to their rigor and precision, formal methods have been utilized for specifying, analyzing, and verifying various artifacts in MDE. The literature is abundant with contributions on the application of formal methods for automating various MDE tasks. However, we limit ourselves to a few examples of their most common uses.
@@ 2.2.2. Search-based Techniques
These techniques have emerged as valuable tools for automating various activities of MDE. This line of research views the automation of many MDE tasks as optimization problems, which can be effectively tackled using search-based methods (Boussaïd et al., 2017). One of the earliest applications of this concept was in the domain of model transformations, particularly using examples of input-output models. Researchers utilized search-based techniques for model diffing based on refactoring detection(Kessentini et al., 2017), model merging (Mansoor et al., 2015), and to perform model transformations (Kessentini et al., 2008) and even to learn transformation rules (Faunes et al., 2013; Baki and Sahraoui, 2016).
@@
Another significant application of search-based techniques in MDE is the repair of modeling artifacts. The evolution of metamodels can invalidate many associated artifacts due to their dependency on the type systems defined within the metamodels. Evolutionary algorithms have been employed to repair models (Kessentini and Alizadeh, 2022), OCL constraints (Batot et al., 2017), and model transformations (Kessentini et al., 2018a). Similarly and independently from the evolution, search-based approaches have been used to repair errors in MDE artifacts such as transformations (VaraminyBahnemiry et al., 2021).
In the realm of design space exploration (DSE), where finding the best design instance for a given problem is challenging, search-based techniques have proven invaluable. DSE, framed as an optimization problem, involves exploring alternative models to identify those that meet specific criteria (Abdeen et al., 2014).
@@ 2.2.4. Artificial Intelligence
Like many other disciplines, AI has been extensively used as an automation technology for MDE. From the early days of MDE until now, rule-based systems have been employed to automate model transformations, utilizing either general-purpose rule languages such as NéOpus (Revault et al., 1995) or specifically dedicated ones like ATL (Jouault et al., 2008). Soon after, the notion of supporting modelers in writing transformation rules emerged. In this regard, machine learning techniques were applied to derive transformation rules from examples using supervised learning (Varró and Balogh, 2007) or unsupervised learning (Dolques et al., 2009).
@@
The remarkable power of LLMs has partially alleviated this problem. Thus, it became possible to automate several modeling tasks with little or no training data. However, the effectiveness of using LLMs for this purpose varies, as Camara et al. (Cámara et al., 2023) highlight. Ben Chaaben et al. explored the application of LLMs for static and dynamic domain-model completion (Chaaben et al., 2023). Model generation with LLMs has also garnered significant interest among researchers, encompassing diverse model types such as goal models (Chen et al., 2023a) and static domain models (Chen et al., 2023b).
@@ 2.3. Mapping between Activities and Enabling Technologies
Testing / Debugging (Popoola et al., 2016; Bagherzadeh et al., 2017)+ (Hilken et al., 2018; Babikian et al., 2022)+ (Walter et al., 2014) (Batot and Sahraoui, 2016; Gómez et al., 2012)+ (Rahimi et al., 2023; López and Cuadrado, 2023)+
Model transformations (Cuadrado et al., 2022; Biermann et al., 2006)+ (Biermann et al., 2008; Gammaitoni and Kelsen, 2019)+ (Roser and Bauer, 2005; Kappel et al., 2006)+ (Eisenberg et al., 2021) (Kessentini et al., 2008; Baki and Sahraoui, 2016)+ (Burgueño et al., 2022; Lano and Xue, 2023)
@@
(Ruscio et al., 2012; Herrmannsdoerfer, 2011)+ (Taentzer et al., 2012; Levendovszky et al., 2014)+ (Kessentini et al., 2018b; Kessentini and Alizadeh, 2022)+
@@
Modeling
Model generation
Model search
(Lucrédio et al., 2008; Basciani et al., 2018b)+
Model completion
(Kuschke et al., 2013; Born et al., 2008)+ (Sen et al., 2010; Steimann and Ulke, 2013)+ (Bislimovska et al., 2014; López and Cuadrado, 2022)+ (Nechypurenko et al., 2007) (Di Rocco et al., 2021) (Chaaben et al., 2023)
Models at runtime
Model classification
(Rubei et al., 2021; Basciani et al., 2016)+ (Nguyen et al., 2019) (Nguyen et al., 2021)
@@
Table 1121212The table is inevitably nonexhaustive and merely intended to reflect the kinds of activities and supporting approaches available. gives a partial summary of how automation-enabling technologies have been used for various MDE tasks. As we can see, a range of techniques, including direct programming, formal methods, and search methodologies have been extensively employed to automate various MDE tasks. To a lesser extent, the MDE community have leveraged, on the one hand, model sets and ontologies and, on the other hand, other AI techniques to automate some distinct activities. On another note, in recent years, there has been a surge in interest and adoption of deep learning-based approaches for MDE, albeit their widespread use has been hindered by the scarcity of large-scale datasets for model training or refinement.
[…]
@@
Looking ahead, LLMs are poised to become a pivotal research domain in automating MDE tasks. However, their application extends beyond mere model generation or completion tasks. There exists a vast scope for creative exploration and utilization of LLMs in diverse MDE activities, heralding a new era of innovation and efficiency in software engineering practices.
@@ 3. Perspectives and Challenges
Management of ever-increasing diversity of languages and tools: MDE has seen a proliferation of domain-specific modeling languages and tools, which has led to a diverse range of technologies. However, this diversity has posed challenges for generic automation as each modeling language has its own syntax, semantics, and transformation rules. Despite efforts to standardize syntax, the MDE field still needs comprehensive standards, especially concerning semantics, transformation rules, and model interchange formats. Moreover, the rise of low-code/no-code platforms makes this issue even more difficult as adherence to existing standards becomes increasingly sporadic. Thus, it is imperative to develop standardized interfaces, transformation mechanisms, and model interchange formats to ensure interoperability and tooling support across different formalisms, languages, and paradigms.
@@
Combination of symbolic and non-symbolic AI: Integrating symbolic AI techniques, such as logic reasoning and rule-based systems, with non-symbolic approaches, such as neural networks and deep learning, offers opportunities for enhancing automation in MDE. By leveraging the complementary strengths of both paradigms, researchers can develop hybrid AI systems that excel in tasks requiring both symbolic reasoning and pattern recognition. For example, combining symbolic reasoning with neural networks enables MDE tools to interpret and manipulate models based on logical constraints while leveraging the scalability and flexibility of neural networks for data-driven tasks or even exploiting generative AI tools to synthesize models that are domain or context-aware.
@@
Emancipation from formal languages with LLMs: Creating domain-specific modeling languages is a crucial undertaking that involves collaboration between different roles, including language and domain experts. This collaborative effort encompasses various steps, including developing concrete syntaxes, the definition of formal semantics, and the construction of supportive modeling environments. However, the language development process can be challenging, often necessitating refinement of language definitions to refine user requirements or address unforeseen ones. LLMs offer a promising option for the early evaluation of languages by specifying their semantics in natural language and facilitating early analysis and stakeholder interactions to validate whether the language aligns with user requirements.
Template-based code generation vs. AI-based code generation: While template-based methods provide transparency and control over the generated code, AI-based approaches offer adaptability and potential for optimizing code generation based on diverse requirements and contexts. Researchers can develop code-generation techniques by exploring hybrid approaches that combine the strengths of both paradigms.
Model management automation with AI approaches: AI-driven techniques can enhance model management processes by improving semantic understanding and context-aware manipulation of models. By leveraging AI for tasks such as model differencing, merging, and versioning, MDE can achieve greater efficiency and reliability in managing evolving model artifacts. For example, AI algorithms can analyze the semantics of models to detect and resolve (or propose how to resolve) conflicts during model merging, thereby reducing manual effort and potential errors in model integration.
@@
Moving from general LLMs to task-specific AI agents: In the traditional context of MDE, a platform typically denotes a particular technology stack or execution environment. However, with the increasing adoption of LLMs, the platform concept expands to encompass a collaborative ecosystem of multiple language models operating together within a multi-agent system. Specifically, we foresee a paradigm shift towards the widespread adoption of task-specific AI agents driven by LLMs. These specialized agents are designed to execute various model management tasks, such as model evolution, comparison, domain modeling, training and test data generation, and model completion (Hong et al., 2023).
@@
Orchestration of automated MDE tasks: The application of modeling concepts and tools involves executing different tasks, including model transformations, validation, and code generation. To this end, despite the availability of domain-specific languages like Wires (Rivera et al., 2009b), there is still the need for higher-level automation strategies that, starting from user requirements, can coordinate low-level automated tasks and that are flexible to accommodate different modeling languages, tools, and environments. Additionally, achieving a seamless integration of automated processes requires addressing compatibility issues and ensuring interoperability between different automation tools and frameworks.
Democratizing the usage of task-specific AI agents: To enable the usage of task-specific AI agents, it is necessary to devise low-code platforms that empower users to leverage task-specific AI agents seamlessly. Such platforms can facilitate broader participation in MDE activities across domains and expertise levels by providing intuitive interfaces for coordinating diverse AI models. Moreover, low-code platforms can serve as orchestrators for deploying and managing AI agents, allowing users to customize and extend automation capabilities according to their specific requirements and preferences.
Combination of physical/biological/social/software aspects in complex systems: Integrating physical, biological, social, and software aspects into the same system presents a multifaceted challenge that requires a comprehensive approach to modeling and abstraction. MDE stands as an ideal integration vehicle for managing these heterogeneous aspects by providing a unified framework for representing and analyzing complex systems. Digital twins, which are virtual replicas of physical entities or systems, offer a powerful abstraction mechanism for simulating and analyzing the behavior of interconnected systems. By leveraging digital twins, MDE can facilitate the design and optimization of integrated systems across diverse domains, ranging from healthcare to sustainable smart cities.
Emergence of non-traditional interface devices: Integrating speech interfaces, gesture recognition, and other non-traditional input modalities in MDE tools can democratize software development by enabling a broader range of stakeholders to participate in modeling activities. By providing intuitive and accessible interfaces, such as voice commands or touch-based interactions, MDE tools can lower the barriers to entry for users with diverse backgrounds and abilities. Moreover, non-traditional interface devices facilitate collaboration and communication among distributed teams, allowing users to interact with models in real time and co-create solutions collaboratively. To this end, interdisciplinary work at the intersection of human-machine interaction and software engineering research fields will be needed (Kashfi et al., 2019).
Feedback-driven MDE: MDE tools and processes will need built-in feedback mechanisms. This feedback could come from engineers, stakeholders or the system itself (i.e., in the form of real-time data) and could be sporadic or continuous. Tools, processes and teams should not disregard feedback. New methodologies to work in a new environment where automated feedback integration are needed, too. A straightforward scenario is Digital Twins. For instance, Digital Twins can be used for product development, where design and engineering teams can investigate a broader range of design possibilities early in the development, saving money, resources and time. In such a scenario, the engineering processes must be adapted to collect and react to the outputs/feedback received by the Digital Twin, and engineering teams must have methodologies put in place.
Automated and semi-automated decision-making in MDE: As software and systems grow in complexity and the demands for faster delivery increase, automating decision-making processes becomes needed to improve efficiency and quality. On a purely technical level, automated decision-making will help choose the best transformation languages, model checkers, AI algorithms, execution models, etc. Furthermore, it would ensure that decisions are made consistently throughout the life of a project and by different teams. It would also prevent human error in repetitive decision-making processes (e.g., tools enhanced with automated support could exploit data and human-defined objectives to make or suggest optimal decisions).
@@
Scarcity of datasets and benchmarks for AI models: The advancement of AI-driven automation in MDE is hindered by the scarcity of high-quality datasets and benchmarks for training and evaluating AI models. Unlike domains like natural language processing or computer vision, where large-scale datasets are readily available, MDE lacks comprehensive datasets that capture a diverse range of modeling tasks and application domains (López et al., 2024). Moreover, concerns regarding data privacy and the ethical use of data further complicate the acquisition and sharing of modeling datasets. Addressing these challenges requires collaboration between researchers, practitioners, and stakeholders to curate datasets, define evaluation metrics, and establish best practices for data collection and sharing in MDE contexts (Rocco et al., 2015).
@@
The advent of LLMs presents new opportunities for dataset creation in various domains. With the ability to generate human-like text and simulate expert knowledge, LLMs offer a cost-effective alternative to traditional human labeling methods. For instance, LLMs can automatically generate labeled datasets for MDE tasks such as model classification or sentiment analysis. Additionally, LLMs can serve as judges to evaluate the quality and accuracy of tasks performed by other LLMs, thereby facilitating continuous improvement and refinement of AI-driven systems.
Better modeling tools: Despite significant advancements in modeling tools, many existing tools in MDE are tailored towards expert users and lack intuitive interfaces for non-specialists. This poses challenges for automation and democratizing the adoption of MDE methodologies across diverse domains and user communities (Ruscio et al., 2022). Improving the accessibility and usability of modeling tools requires incorporating user-centered design principles, conducting usability studies, and providing comprehensive documentation and training materials. Additionally, ensuring compatibility and interoperability between different modeling tools and environments is essential for facilitating seamless collaboration and exchange of models among users with varying expertise levels.
Teaching people how to prompt: Despite the advancements in AI and automation, human expertise remains indispensable in designing complex systems and formulating effective prompts for AI models. Teaching people how to prompt effectively involves imparting knowledge and skills in understanding system requirements, defining clear objectives, and structuring prompts to elicit desired responses from AI models. Thus, human expertise is essential for ensuring the accuracy, relevance, and ethical considerations of AI-generated outputs.

### v2

@@ 2.2.1. Formal Methods
Due to their rigor and precision, formal methods have been utilized to specify, analyze, and verify various MDE artifacts. The literature is abundant with contributions on the application of formal methods for automating various MDE tasks. However, we limit ourselves to a few examples of their most common uses.
@@
Although formal methods have contributed to explicitly modeling automation problems and provided effective solutions for some MDE task automation, their use is limited by several issues. The most well-known challenge is scalability. Formal methods often struggle when applied to large and complex models due to the computational cost of techniques like model checking and constraint solving. Moreover, certain MDE tasks, such as creative or exploratory modeling, are not easily automatable using solely formal methods. Finally, despite their potential, tools supporting formal methods often lack a direct integration with mainstream MDE environments, which hampers seamless workflows.
@@ 2.2.2. Search-based Techniques
These techniques have emerged as valuable tools for automating various MDE activities. This line of research views the automation of many MDE tasks as optimization problems, which can be effectively tackled using search-based methods (Boussaïd et al., 2017). One of the earliest applications of this concept was in the domain of model transformations, particularly using examples of input-output models. Researchers utilized search-based techniques for model diffing based on refactoring detection(Kessentini et al., 2017), model merging (Mansoor et al., 2015), and to perform model transformations (Kessentini et al., 2008) and even to learn transformation rules (Faunes et al., 2013; Baki and Sahraoui, 2016).
@@
Another significant application of search-based techniques in MDE is repairing modeling artifacts. The evolution of metamodels can invalidate many associated artifacts due to their dependency on the type systems defined within the metamodels. Evolutionary algorithms have been employed to repair models (Kessentini and Alizadeh, 2022), OCL constraints (Batot et al., 2017), and model transformations (Kessentini et al., 2018a). Similarly and independently from the evolution, search-based approaches have been used to repair errors in MDE artifacts such as transformations (VaraminyBahnemiry et al., 2021).
In the realm of design space exploration (DSE), where finding the best design instance for a given problem is challenging, search-based techniques have shown invaluable capabilities. DSE, framed as an optimization problem, involves exploring alternative models to identify those that meet specific criteria (Abdeen et al., 2014).
@@
In summary, search-based techniques have facilitated the automation of various MDE tasks by offering algorithms not only to solve specific instances but also to learn actionable knowledge that can be leveraged to streamline these tasks. However, this family of techniques has also several limitations. Tasks involving large search spaces are challenging to manage and often result in suboptimal solutions, particularly when determining the appropriate search parameters is difficult. Furthermore, the effectiveness of search-based techniques heavily relies on the quality of the objective functions guiding the search. Defining suitable, measurable, and well-balanced objectives within the context of MDE can be a significant challenge. Lastly, the solutions generated by search-based techniques may overfit specific problem instances, reducing their reusability and limiting their generalization to other MDE scenarios.
@@ 2.2.3. Extensional and Intensional Knowledge
Capturing and formalizing the knowledge required for automation, whether as artifacts or ontologies, is a complex and resource-intensive task. Automation solutions that rely on both types of knowledge are often highly domain-specific, limiting their generalizability and requiring significant effort to adapt the knowledge for use in new domains. The effectiveness of extensional knowledge depends greatly on the quality, quantity, and relevance of the artifacts in the repository. Currently, existing repositories are often restricted to specific formalisms, such as UML class diagrams, making it challenging to assess the quality and representativeness of the artifacts. Additionally, maintaining large collections of models or transformations can become burdensome, with ensuring that repositories remain consistent and up-to-date posing a significant challenge. Finally, although ontologies provide semantic richness, designing and maintaining them for complex domains is both time-intensive and requires specialized expertise in knowledge engineering.
@@ 2.2.4. Artificial Intelligence
Like many other disciplines, artificial intelligence has been extensively used as an automation technology for MDE. From the early days of MDE until now, rule-based systems have been employed to automate model transformations, utilizing either general-purpose rule languages such as NéOpus (Revault et al., 1995) or specifically dedicated ones like ATL (Jouault et al., 2008). Soon after, the notion of supporting modelers in writing transformation rules emerged. Machine learning techniques were applied to derive transformation rules from examples using supervised learning (Varró and Balogh, 2007) or unsupervised learning (Dolques et al., 2009).
@@
The remarkable power of Large Language Models (LLMs) [^17] Note that a model in AI and a software model do not refer to the same type of model. In the AI field, a model is the result of analyzing datasets to find patterns and make predictions (e.g., it could be a trained neural network or a statistical model). To avoid confusion, in this work, each time we refer to an AI model, we always refer to it with a precise name (e.g., AI model, ML model, Language model, …) and never as “model” alone. When we use “model” alone, it always refers to a software model. has partially alleviated this problem. Thus, automating several modeling tasks with little or no training data became possible. However, the effectiveness of using LLMs for this purpose varies, as Camara et al. (Cámara et al., 2023) highlight. Ben Chaaben et al. explored the application of LLMs for static and dynamic domain-model completion (Chaaben et al., 2023). Model generation with LLMs has also garnered significant interest among researchers, encompassing diverse model types such as goal models (Chen et al., 2023a) and static domain models (Chen et al., 2023b).
Unlike other fields, the MDE research community has yet to fully capitalize on the potential of AI and, in particular, deep learning (DL) and large language models (LLMs) for automation, largely due to several inherent challenges. One major obstacle is the scarcity, fragmentation, and inconsistency of MDE-specific datasets, which are essential for effectively training and fine-tuning DL models. Moreover, MDE involves diverse, highly structured artifacts closely tied to domain-specific languages (DSLs). Adapting DL architectures to accommodate this diversity and to work seamlessly with domain-specific representations remains an unresolved challenge. Another significant obstacle in leveraging LLMs for MDE tasks lies in the challenge of crafting effective prompts to extract relevant knowledge from the LLMs and mapping this knowledge accurately to the DSLs in which it must be represented. This complexity is compounded by the reliance of MDE practitioners on well-established toolchains. Integrating DL and LLM-based solutions into these ecosystems is challenging, particularly when existing tools are not designed to address the unique demands of MDE-specific tasks.
@@ 2.3. Mapping between Activities and Enabling Technologies
Modeling
Model generation
Model search
(Lucrédio et al., 2008; Basciani et al., 2018b)+ (López and Cuadrado, 2022)
Model completion
(Kuschke et al., 2013; Born et al., 2008)+ (Sen et al., 2010; Steimann and Ulke, 2013)+ (Bislimovska et al., 2014; Burgueño et al., 2021)+ (Nechypurenko et al., 2007) (Di Rocco et al., 2021) (Chaaben et al., 2023; Tinnes et al., 2024)
Models at runtime
Model classification
(Rubei et al., 2021; Basciani et al., 2016)+ (Nguyen et al., 2019; López et al., 2022) (Nguyen et al., 2021)
@@
Testing / Debugging (Popoola et al., 2016; Bagherzadeh et al., 2017)+ (Hilken et al., 2018; Babikian et al., 2022)+ (Walter et al., 2014) (Batot and Sahraoui, 2016; Gómez et al., 2012)+ (Rahimi et al., 2023; López and Cuadrado, 2023)+ (Cámara et al., 2024)
Model transformations (Cuadrado et al., 2022; Biermann et al., 2006)+ (Biermann et al., 2008; Gammaitoni and Kelsen, 2019)+ (Roser and Bauer, 2005; Kappel et al., 2006)+ (Eisenberg et al., 2021) (Kessentini et al., 2008; Baki and Sahraoui, 2016)+ (Burgueño et al., 2022; Lano and Xue, 2023) (Kazai and Osei, 2024; Buchmann, 2024)
@@
(Ruscio et al., 2012; Herrmannsdoerfer, 2011)+ (Taentzer et al., 2012; Levendovszky et al., 2014)+ (Kessentini et al., 2018b; Kessentini and Alizadeh, 2022)+ (Kebaili et al., 2024)
@@
Table 1 gives a summary of how automation-enabling technologies have been used for various MDE tasks. [^19] Table 1 is inevitably non-exhaustive and merely intended to reflect the kinds of activities and available supporting approaches. Many techniques, including direct programming, formal methods, and search methodologies have been extensively employed to automate various MDE tasks. To a lesser extent, the MDE community has leveraged, on the one hand, model sets, and ontologies and, on the other hand, other AI techniques to automate some distinct activities. On another note, in recent years, there has been a surge in interest and adoption of deep learning-based approaches for MDE. However, the scarcity of large-scale datasets for model training or refinement has hindered their widespread use.
The utilization of Language Model-based systems and other cutting-edge technologies in this realm is still in its nascent stages. To date, the primary focus has been on model generation and completion tasks within MDE workflows. Nonetheless, this landscape presents numerous untapped research directions for further exploration and refinement of deep learning approaches in MDE.
@@
Looking ahead, LLMs are poised to become a pivotal research domain in automating MDE tasks. However, their application extends beyond mere model generation or completion tasks. A vast scope exists for creative exploration and utilization of LLMs in diverse MDE activities, heralding a new era of innovation and efficiency in software engineering practices.
@@ 3. Perspectives and Challenges
3.1. AI and MDE
As with many disciplines, AI significantly enhances the potential for automating tasks. In the past decade, considerable advances have been made, particularly in natural language and code models. MDE is ideally positioned to benefit from both types of LLMs, as it uniquely represents software at a high level of abstraction while also translating these abstractions into operational implementations. However, these opportunities come with many challenges, as AI models have not been specifically designed for these uses, unlike code models, whose design is aligned with coding tasks. In this section, we discuss these opportunities and challenges.
@@ 3.1. AI and MDE
Combination of symbolic and non-symbolic AI: Integrating symbolic AI techniques—such as logic reasoning and rule-based systems—with non-symbolic approaches like neural networks and deep learning presents significant opportunities for advancing automation in MDE. By harnessing the complementary strengths of these paradigms, researchers can develop hybrid AI systems that excel in tasks requiring both symbolic reasoning and pattern recognition. For instance, combining symbolic reasoning with neural networks allows MDE tools to interpret and manipulate models according to logical constraints, while leveraging the scalability and adaptability of neural networks for data-driven tasks. Additionally, generative AI tools can synthesize domain-specific or context-aware models, further enhancing the effectiveness and applicability of MDE tools.
@@
Emancipation from formal languages with LLMs: Creating domain-specific modeling languages is a critical endeavor that requires collaboration among various roles, including language and domain experts. This process involves several key steps: developing concrete syntaxes, defining formal semantics, and constructing supportive modeling environments. However, the language development process often presents challenges, such as refining language definitions to meet evolving user requirements or address unforeseen issues. Additionally, the resulting languages may not always be user-friendly for domain experts. LLMs offer a promising solution for the early evaluation of these languages. By specifying their semantics in natural language, LLMs can facilitate preliminary analysis and stakeholder interactions, helping to ensure that the language aligns with user needs. This approach can be further extended by using natural language to represent different aspects of a domain and allowing LLMs to manage ambiguity in these representations for subsequent development tasks.
@@
Moving from general LLMs to task-specific AI agents: In the traditional MDE context, a platform typically refers to a specific technology stack or execution environment. However, with the growing adoption of LLMs, the concept of a platform broadens to encompass a collaborative ecosystem of multiple language models operating within a multi-agent system. We anticipate a paradigm shift toward the widespread use of task-specific AI agents powered by LLMs. These specialized agents are designed to handle various model management tasks, including model evolution, comparison, domain modeling, training and test data generation, and model completion (Hong et al., 2023).
@@
Orchestration of automated MDE tasks: Building on the previous point, applying modeling concepts and tools involves performing various tasks, such as model transformations, validation, and code generation. Despite the availability of domain-specific languages like Wires (Rivera et al., 2009b), there remains a need for higher-level automation strategies that can, starting from user requirements, orchestrate low-level automated tasks while remaining flexible enough to support different modeling languages, tools, and environments. Furthermore, achieving seamless integration of these automated processes requires overcoming compatibility issues and ensuring interoperability between diverse automation tools and frameworks.
Model management automation with AI approaches: AI-driven techniques can significantly enhance model management processes by improving the semantic understanding and context-aware manipulation of models. By applying AI to tasks such as model differencing, merging, and versioning, MDE can achieve higher efficiency and reliability in handling evolving model artifacts. For instance, AI algorithms can analyze the semantics of models to detect and resolve—or propose solutions for—conflicts during model merging, reducing manual effort and minimizing errors in model integration.
AI to improve reuse in MDE: Unlike in the realm of code, the level of artifact reuse in MDE remains low. This can be partly attributed to the limited presence of explicit semantics associated with these artifacts and the vast and ambiguous vocabulary often employed. However, the ability of LLMs to grasp the meaning of representations unlocks new possibilities for reuse. This includes enhanced model search, model completion, and even the generation of model transformations.
Template-based code generation combined with AI-based code generation: While template-based methods provide transparency and control over the generated code, AI-based approaches offer adaptability and potential for optimizing code generation based on diverse requirements and contexts (Weyssow et al., 2024). Researchers can develop code generation techniques by exploring hybrid approaches that combine the strengths of both paradigms. Also, for emerging platforms, an important research challenge is how code generation facilities can be provided as reference applications from which code generation schemes may be abstracted. This challenges both paradigms, which may be mitigated by having hybrid code generators that can resort to human-defined templates combined with AI-based code injections to balance the required human efforts as well as the demand for large sets of reference applications.
@@
Scarcity of datasets and benchmarks: The scarcity of high-quality datasets hinders the advancement of AI-driven automation in MDE and benchmarks for training and evaluating AI models. Unlike domains such as natural language processing or computer vision, where large-scale datasets are readily available, MDE lacks comprehensive datasets that capture a diverse range of modeling tasks and application domains (López et al., 2024). In addition, concerns about data privacy and the ethical use of data further complicate the acquisition and sharing of modeling datasets. Addressing these challenges requires collaboration between researchers, practitioners, and stakeholders to curate datasets, define evaluation metrics, and establish best practices for data collection and sharing in MDE contexts (Rocco et al., 2015; Cámara et al., 2024).
@@
The advent of LLMs presents new opportunities for the creation of datasets in various domains. With the ability to generate human-like text and simulate expert knowledge, LLMs offer a cost-effective alternative to traditional human labeling methods. For instance, LLMs can automatically generate labeled datasets for MDE tasks such as model classification or sentiment analysis. Additionally, LLMs can serve as judges to evaluate the quality and accuracy of tasks performed by other LLMs, thereby facilitating continuous improvement and refinement of AI-driven systems.
3.2. Human Factors and MDE
Models enable communication among stakeholders, software engineers, and final users. Models and MDE also increase the productivity of software development teams thanks to the (partial) automation of the development process (Brambilla et al., 2017). When using models, humans have goals and expectations, which should be met as both the success of their projects as well as the modeling disciple depend on this. Therefore, special attention should be paid to how humans understand and interact with models. As MDE processes are being automated and the field is evolving, exploration of how these changes affect humans needs to be explored. It is also worth noting that Liebel et al. have reported that “there is an underrepresentation of research focusing on human factors in modeling” (Liebel et al., 2024). In the following, we discuss the opportunities and challenges that the latest advances in MDE are bringing to the field.